### v1

# High-throughput electro-optic upconversion and downconversion with few-photon added noise

Abstract
A microwave-optical transducer of sufficiently low noise and high signal transfer rate would allow entanglement to be distributed between superconducting quantum processors at a rate faster than the lifetimes of the quantum memories being linked. Here we present measurements of a membrane-based opto-electromechanical transducer with high signal throughput, as quantified by an efficiency-bandwidth-duty-cycle product of 7 kHz, approaching quantum-enabled operation in upconversion as well as downconversion, with input-referred added noise of 3 photons. In downconversion, throughput of this magnitude at the few-photon noise level is unprecedented. Using the quantum channel capacity, we also find an expression for the maximum rate at which quantum information can be transduced, providing insight into the importance of improving both a transducer’s throughput and noise performance. With feasible improvements, the high throughput achieved with this device positions membrane-based transducers as a strategic choice for demonstrations of a quantum network with reasonable averaging times.
A transducer bridging microwave and optical frequencies with sufficiently low noise would allow the deterministic generation of arbitrary optical quantum states for applications in metrology, the storage and manipulation of optical quantum information, and the networking of remote superconducting quantum processors. Reducing input-referred added noise $N_{\text{add}}$ to as low a level as possible is imperative for such quantum state manipulation. At the same time, it is useful to increase the rate of transduction to enable transducing signals at rates that are compatible with the lifetimes of logical qubits that process the transduced quantum information. Current state-of-the-art lifetimes in superconducting architectures are approximately 1 ms using physical qubits [1, 2], 1 ms using error-corrected bosonic encodings [3], and 300 $\mu$s using logical encodings of many physical qubits [4]. Finite qubit lifetimes place constraints not only on the transducer bandwidth $B$ (which must be fast enough to transduce quantum information within the time qubits can faithfully store it), but also on the overall signal throughput in order to support an entangled bit rate much faster than decoherence rates. Signal throughput is further reduced by less-than-unity transducer efficiency $\eta$ and duty cycle $D$, underscoring the importance of increasing the throughput of quantum transducers, even as progress is made on logical lifetimes of quantum memories. In practice, there is often a tradeoff between noise performance and throughput, as increasing throughput frequently involves increasing the average drive power, which can lead to heating-induced noise.
Concurrent with improving the performance of electro-optic transducers, a fully developed transducer platform would be capable of operating in both directions of frequency conversion. For instance, there are several proposed protocols for sharing entanglement between separated quantum systems, and these place different demands on the transducers in a network of microwave-frequency nodes linked optically. Long-range quantum networks dominated by loss will likely benefit from using heralding protocols, in which entanglement infidelity due to inefficiency is mitigated at the cost of transmission rates and greater noise sensitivity [5]. Such a protocol favors a network topology with transducers operating in upconversion (from microwave to optical frequencies) to take advantage of mature optical single-photon-detection technologies [6]. On the other hand, transducers operating in downconversion (from optical to microwave frequencies) would be useful for generating entanglement from squeezed optical light [7]. Further, deterministic state transfer protocols [8, 9] are possible if efficiencies are high, and require one transducer operating in upconversion and one in downconversion. In general, a transducer’s performance can be asymmetric between the two directions of frequency conversion, depending on the origin of the dominant source of noise. Published work to date has typically operated more favorably in upconversion than in downconversion, in the relatively few works where downconversion performance has been demonstrated and quantified [10, 11].
In this work, we present measurements of a transducer with high throughput, $\Theta=\eta BD$ exceeding 7 kHz, and noise performance close to input-referred added noise $N_{\text{add}}=1$, expressed in units of photons, or photons per second per Hz of bandwidth, in upconversion ($N_{\text{add}}^{\text{up}}=2.6$) as well as downconversion ($N_{\text{add}}^{\text{down}}=3$). While similar throughputs have been achieved in a device with quantum-enabled performance in upconversion [12], in downconversion, our device increases throughput by three orders of magnitude over the leading quantum demonstration [11], at the cost of a single order of magnitude worse noise performance. Our platform achieves such performance due to its robustness to strong optical pumping, as demonstrated in previous work [13], as well as significant improvement to the electrical circuit that we show here.
As transducers push the boundaries of both lower noise and higher throughput, it is important to understand the impacts of each of these two parameters in quantum networks, beyond the requirement $N_{\text{add}}<1$. We approach this analysis by mathematically integrating the quantum channel capacity, in the case of transducers with spectrally flat, thermal input-referred added noise, and find an expression that assesses transducer utility using noise and throughput.
Figure 1: Upconversion transducer performance with $B=22$ kHz. The apparent efficiency $\eta$ quantifies the fraction of the incident microwave signal (blue circles) that is transduced to optical frequencies (red circles), rather than lost (empty dashed circles). The upconversion input-referred added noise $N_{\text{add}}^{\text{up}}$ (gray dashed circles) compares the noise rate measured at the output $N_{\text{out}}^{\text{up}}$ (gray circles) to a potential signal at the input. (a) Apparent transduction efficiency $\eta$ vs. signal detuning from pump $\omega/2\pi$. (b) Output-referred upconversion noise spectrum. (c) Input-referred added noise spectrum in upconversion. The grayed-out regions indicate frequency bands contaminated by noise from the high thermal occupation of optomechanically stiff modes. Excluding these regions and integrating the noise across the bandwidth of the transducer yields $\bar{N}^{\text{up}}_{\text{add}}=2.7$ (horizontal dashed line).
Figure 2: Downconversion transducer performance with $B=19$ kHz. (a) Apparent transduction efficiency $\eta$ vs. signal detuning from pump $\omega/2\pi$. (b) Downconversion output-referred noise spectrum. The broadband noise contribution extending beyond the damped linewidth of the transduction mode is due to unwanted additional occupation of the LC circuit resulting from the microwave pump. Destructive interference between this microwave noise and the mechanical motion it drives results in noise squashing (Eq. 2). (c) Input-referred added noise spectrum in downconversion. Broadband circuit noise divided by the frequency-dependent $\eta$ results in $N_{\text{add}}^{\text{down}}$ increasing with detuning from the center of the transduction bandwidth. Additionally, noise squashing near resonance suppresses the noise level near $\omega_{\text{m}}/2\pi=1.27$ MHz.
.1 Characterizing high-throughput performance
Our transducer is a doubly-parametric membrane-based transducer of similar design to Refs. [13, 14], with the following improvements: the silicon nitride Si3N4 in the device is annealed [15], and the membrane is “pinned” in place with a silicon post to more reproducibly set the transducer’s capacitor spacing [16, 17]. Furthermore, we have improved the noise performance of the optical and microwave pumps with additional filtering (Methods). The mechanical mode of the Si3N4 membrane with resonant frequency $\omega_{\text{m}}=2\pi\cdot 1.27$ MHz couples an optical Fabry-Pérot cavity and a microwave-frequency superconducting LC circuit via the optomechanical and electromechanical interactions [18]. Electrical and optical pumps red-detuned by $\omega_{\text{m}}$ from their respective resonators enhance these interactions, coupling the mechanical mode to the microwave (optical) field at rate $\Gamma_{\text{e}}$ ($\Gamma_{\text{o}}$), proportional to the pump power. Imperfect suppression of the Stoke’s sideband of the transduction mode leads to gain $\mathcal{A}=\mathcal{A}_{\text{e}}\mathcal{A}_{\text{o}}$, where $\mathcal{A}_{\text{e}}$ ($\mathcal{A}_{\text{o}}$) is the portion of the gain due to the electromechanical (optomechanical) interaction, which is undesirable because it is accompanied by additional noise [18, 19]. The apparent efficiency $\eta=\mathcal{A}\eta_{\text{t}}$ includes this gain factor and gives the fraction of the incident signal that exits the transducer, where $\eta_{\text{t}}=\eta_{\text{M}}(4\Gamma_{\text{e}}\Gamma_{\text{o}}/\Gamma_{\text{T}}^{2})$ is the ideal transduction efficiency of a device operating with perfect optomechanical and electromechanical sideband resolution [13], and the maximum achievable efficiency $\eta_{\text{M}}$ is limited by loss in the electromagnetic resonators and an optical mode matching factor. Our platform achieves the highest external microwave-optical transduction efficiencies demonstrated to date [13]. In this device we achieve apparent efficiencies of $\eta=0.4$, calibrated by the procedure used in Ref. [20]. The transduction bandwidth $B=\Gamma_{\text{T}}/2\pi$ is determined by the total loss rate of the mechanical mode, including the pump-enhanced loss rates, $\Gamma_{\text{T}}=\Gamma_{\text{e}}+\Gamma_{\text{o}}+\gamma_{\text{m}}$, where $\gamma_{\text{m}}\ll\Gamma_{\text{e}},\Gamma_{\text{o}}$ is the intrinsic loss rate of the transduction mode.
The measured bidirectional apparent efficiency $\eta(\omega)=\sqrt{\epsilon_{\text{PL}}}|S_{\text{eo}}(\omega)||S_{\text{oe}}(\omega)|/|S_{\text{ee}}(\omega)||S_{\text{oo}}(\omega)|$ is shown in Fig. 1(a), where e.g. $S_{\text{oe}}$ denotes the scattering parameter governing a signal incident on the port electrical port “e” and detected from the optical port “o”, and $\epsilon_{\text{PL}}$ is a factor quantifying the spatial mode matching of the optical pump and local oscillator beam used for heterodyne readout [18]. The efficiency follows a Lorentzian lineshape corresponding to the susceptibility of the damped mechanical transduction mode. In addition to the optomechanically compliant mode that we use for transduction, our device also exhibits optomechanically stiff modes, characterized by significant motion in the massive substrate that makes them resistant to optomechanical damping. The dip in the efficiency at frequency $\omega_{\text{s}}/2\pi=1.275$ MHz and approximate width of 1 kHz in Fig. 1(a) is due to one such mode, and is the result of destructive interference of the coherent signal tone propagating through both the stiff mode and the compliant mode. The prominence of stiff modes in this generation of device is due to an error in the fabrication geometry, which can be addressed in subsequent designs (Methods).
The output-referred noise spectrum $N_{\text{out}}^{\text{up}}=N_{\text{det,o}}/\xi_{\text{o}}$ (Fig. 1(b)) corresponds to the noise measured at the optical heterodyne detection $N_{\text{det,o}}$ in units of photons/s/Hz, divided by the total detection efficiency $\xi_{\text{o}}$ between the output of the transducer and the detected voltage (see Ref. [13] for additional details). Here we measure $\xi_{\text{o}}$ by calibrating the optomechanically measured spectrum in phonon units using sideband asymmetry thermometry. In addition to the thermomechanical noise following the Lorentzian profile of the damped compliant mode, the spectrum shows structure due to multiple stiff modes whose large thermally driven motion is written onto the optically measured output spectrum. The Lorentzian fit (dashed curve) excludes the stiff modes, to provide a guide to the eye. The fit departs from the measurement in the tails because of a small asymmetry in the spectrum due to interference with a highly damped membrane mode with a 950 kHz resonant frequency.
Whereas previous work has focused on the noise level at the center of the transduction bandwidth, in order to more transparently assess the potential of transducing a signal, we compute the frequency-dependent input-referred added noise $N_{\text{add}}^{\text{up}}$ (Fig. 1(c)) by referring the optical output spectrum shown in Fig. 1(b) to the input of the device by dividing by an interpolation of the frequency-dependent transduction efficiency in Fig. 1(a): $N_{\text{add}}^{\text{up}}(\omega)=N_{\text{out}}^{\text{up}}(\omega)/\eta(\omega)$. Neglecting the features due to the stiff modes, the $N_{\text{add}}^{\text{up}}$ spectrum is relatively flat over the frequency band relevant for the transducer. Notching out the frequencies corresponding to the grey regions in Fig. 1(c) and integrating the total contributed noise, we obtain an average added noise of $\bar{N}^{\text{up}}_{\text{add}}=2.7(3)$ over the bandwidth of the transducer (Methods), close to the value of the input referred added noise at the transducer’s maximum efficiency, as determined by the fit to the noise peak, $N_{\text{add}}^{\text{up}}(\omega_{m})=2.6(3)$. The uncertainty in both $\bar{N}^{\text{up}}_{\text{add}}$ and $N_{\text{add}}^{\text{up}}(\omega_{\text{m}})$ is dominated by the statistical uncertainty of the fit parameters in the sideband asymmetry thermometry measurement used to extract $\xi_{\text{o}}$ and calibrate this $N_{\text{add}}^{\text{up}}(\omega)$ spectrum. If we were to physically implement such a sharp filter in a quantum network, the ringing of the narrow filter would only modestly impact the throughput and noise performance (Methods). Depending on the application, shaping the signal pulses to avoid problematic frequency regions or using a selective demodulation envelope in detection are also viable solutions to such frequency-dependent noise profiles. Of course, the more practical solution to the additional noise of the stiff modes is to eliminate them altogether.
The operating parameters used in Fig. 1 achieve $B=22$ kHz, which represent an improvement in bandwidth by approximately two orders of magnitude compared with the device measured in Ref. [13]. Though it may be expected that such an increase in optomechanical and electromechanical damping rates would come at the cost of a commensurate increase in the noise, we measure similar baseline noise levels, i.e. neglecting the noise caused by the stiff modes. This improvement in the robustness of our transducer to strong driving is due to improvement of the single-photon electromechanical coupling rate $g_{\text{e}}$, reduction in phase noise of both optical and microwave pumps from additional filtering (Methods), and annealing the Si3N4 to remove two-level-system defects [15] that were a source of internal loss and, as we find here (Methods), noise in the microwave circuit.
Fig. 2 shows downconversion measurements analogous to those in Fig. 1. Unlike in upconversion, wideband noise extends beyond the bandwidth of the transduction mode due to occupation of the microwave resonator (Fig. 2(b)). This additional noise is associated only with driving the circuit strongly with the microwave pump; we do not observe additional occupation associated with the strong optical pump. Because the microwave noise field also drives the mechanical mode, there is interferometric cancellation at the mechanical frequency leading to squashing measured at the output port [21, 22]. The deficit of noise seen at this output is instead routed out the optical port optomechanically. In Fig. 2(c), in addition to $N_{\text{add}}^{\text{down}}$ being suppressed from squashing near mechanical resonance, $N_{\text{add}}^{\text{down}}$ increases with further detuning from the transduction mode because the relatively broadband noise from microwave circuit occupation is divided by the frequency-dependent conversion efficiency. Unlike upconversion performance, the $N_{\text{add}}^{\text{down}}(\omega)$ spectrum is frequency dependent, and the impact of the off-resonant noise would depend on the details of the network the transducer is embedded in and the protocol used. We therefore for simplicity report downconversion performance using the noise at peak efficiency $N_{\text{add}}^{\text{down}}(\omega_{\text{m}})$ (horizontal dashed line), determined from the minimum of the fit to $N_{\text{out}}^{\text{down}}(\omega)$ referred to the optical input (dashed curve), in line with previous work [20, 23, 13], and analogous to applying additional filtering on $N_{\text{out}}^{\text{down}}$ [10]. The increase in statistical noise in the spectrum measurements shown in Fig. 2 relative to those in Fig. 1 is due to the difference between the optical measurement efficiency $\xi_{\text{o}}=0.4$ and the measurement efficiency of the HEMT amplifier and microwave detection chain $\xi_{\text{e}}=0.01$. The reduced signal-to-noise ratio made calibrating $\xi_{\text{e}}$ electromechanically via sideband asymmetry impractical, so instead we took advantage of our simultaneous optomechanical and electromechanical measurement of the prominent stiff modes (Methods).
.2 Contributions to added noise
To explore noise performance and its differences in upconversion and downconversion, we distinguish expressions for $N_{\text{add}}^{\text{up}}(\omega_{\text{m}})$ and $N_{\text{add}}^{\text{down}}(\omega_{\text{m}})$, making the simplifying assumptions of lossless cavities operating in the resolved sideband regime $\kappa_{\text{e}},\kappa_{\text{o}}\ll 4\omega_{\text{m}}$ with perfect optical mode matching:
$N_{\text{add}}^{\text{up}}(\omega_{\text{m}})$ $=\underbrace{\frac{n_{\text{th}}\gamma_{\text{m}}}{\Gamma_{\text{e}}}+n_{\text{em}}+\frac{n_{\text{om}}\Gamma_{\text{o}}}{\Gamma_{\text{e}}}}_{N_{\text{add,m}}}+\underbrace{\frac{\bar{n}_{\text{o}}\Gamma_{\text{T}}^{2}}{\Gamma_{\text{e}}\Gamma_{\text{o}}}}_{N_{\text{add,e\kern-0.50003pt/\kern-0.74997pto}}}-\underbrace{\frac{2\bar{n}_{\text{o}}\Gamma_{\text{T}}}{\Gamma_{\text{e}}}}_{N_{\text{add,corr}}}$ (1)
and
$N_{\text{add}}^{\text{down}}(\omega_{\text{m}})$ $=\underbrace{\frac{n_{\text{th}}\gamma_{\text{m}}}{\Gamma_{\text{o}}}+n_{\text{om}}+\frac{n_{\text{em}}\Gamma_{\text{e}}}{\Gamma_{\text{o}}}}_{N_{\text{add,m}}}+\underbrace{\frac{\bar{n}_{\text{e}}\Gamma_{\text{T}}^{2}}{\Gamma_{\text{o}}\Gamma_{\text{e}}}}_{N_{\text{add,e\kern-0.50003pt/\kern-0.74997pto}}}-\underbrace{\frac{2\bar{n}_{\text{e}}\Gamma_{\text{T}}}{\Gamma_{\text{o}}}}_{N_{\text{add,corr}}},$ (2)
where $n_{\text{th}}$ is the environmental mechanical occupation, $\bar{n}_{\text{o}}$ is the effective occupation of the optical cavity due to thermal occupation or the presence of technical noise on the optical pump, and $n_{\text{om}}=\bar{n}_{\text{o}}+n_{\text{min,o}}$ is the bath occupation that the mechanical mode is coupled to optomechanically, with $n_{\text{min,o}}$ denoting the optomechanical backaction limit [24]. The quantities associated with the electrical port $n_{\text{em}}$, $\bar{n}_{\text{e}}$, and $n_{\text{min,e}}$ are defined analogously. In the resolved-sideband limit, $n_{\text{om}}=\bar{n}_{\text{o}}$, but we differentiate these noise contributions explicitly here to clarify their physical origins. In Eqs. 1 and 2, the first three terms comprise noise from motion of the transduction mode $N_{\text{add,m}}$, and the fourth term $N_{\text{add,e\kern-0.70004pt/\kern-1.04996pto}}$ is from noise purely of electromagnetic origin measured at the output port. The fifth term $N_{\text{add,corr}}$ results from destructive interference between the fields responsible for the third and fourth terms.
Noise performance between upconversion and downconversion can be asymmetric. In the limit $\Gamma_{\text{o}}=\Gamma_{\text{e}}$, the impedence matching criterion that achieves unit efficiency, Eqs. 1 and 2 become the same (for lossless cavities), even in the presence of additional electromagnetic noise at the output port. However, when $n_{\text{em}}$, $n_{\text{om}}$, and $n_{\text{th}}$ differ, often better performance in one direction can be achieved by mismatching the pump strengths. Practically, the most significant physical effect causing asymmetric transducer performance is that occupations are in general pump-power dependent quantities, and therefore can depend on $\Gamma_{\text{e}}$ and $\Gamma_{\text{o}}$:
$n_{\text{th}}(\Gamma_{\text{e}},\Gamma_{\text{o}});\bar{n}_{\text{e}}(\Gamma_{\text{e}},\Gamma_{\text{o}});\bar{n}_{\text{o}}(\Gamma_{\text{e}},\Gamma_{\text{o}}).$ (3)
In particular, transduction experiments have observed occupations of the form $n_{\text{th}}(\Gamma_{\text{o}})$ [12, 11], $\bar{n}_{\text{e}}(\Gamma_{\text{e}})$ [13], or $\bar{n}_{\text{e}}(\Gamma_{\text{o}})$ [11]. Terms of the form $\bar{n}_{\text{e}}(\Gamma_{\text{o}})$ or $n_{\text{th}}(\Gamma_{\text{o}})$ are common because of the high energy of optical photons compared with the lower frequency scales of the microwave and mechanical excitations used in transducers, and are ruinous for downconversion performance at high optical power and high duty cycle. Though Eqs. 1 and 2 apply specifically to doubly-parametric transducers [13, 12], the noise in other architectures also depends on the transducer’s thermal environment and coupling rates, and similar expressions can be written. The fact that our membrane-based platform does not exhibit noise terms proportional to $\Gamma_{\text{o}}$ is the reason for its downconversion noise performance, even though we operate it with high optical circulating powers of approximately 3 mW in this work.
In upconversion, because of its low optical noise and sufficient sideband resolution, our transducer’s noise performance is determined by the first two terms in Eq. 1 (i.e. $n_{\text{th}}\gamma_{\text{m}}/\Gamma_{\text{e}}+n_{\text{em}}$), as long as $\Gamma_{\text{e}}$ and $\Gamma_{\text{o}}$ are reasonably well-matched. Notably, this approximation implies that $N_{\text{add}}^{\text{up}}$ is independent of $\Gamma_{\text{o}}$. As we observe the microwave circuit occupation to follow the form $\bar{n}_{\text{e}}=a_{\text{e}}\Gamma_{\text{e}}+b_{\text{e}}$ [13] (Methods), the contributions to $N_{\text{add}}^{\text{up}}$ consist of a component of thermomechanical origin that scales as $1/\Gamma_{\text{e}}$, and a component due to LC circuit occupation that increases linearly with $\Gamma_{\text{e}}$. Optimal noise performance in upconversion is achieved by tuning $\Gamma_{\text{e}}$ to balance these two contributions (see Fig. 3(a) below).
Figure 3: Transducer performance. Device performance in upconversion (a) and downconversion (b). Maroon diamonds are measurements of our transducer taken while sweeping $\Gamma_{\text{e}}$ with fixed $\Gamma_{\text{o}}=2\pi\cdot 11$ kHz in (a), and $\Gamma_{\text{o}}/2\pi$ roughly fixed in (b), in the range 9-12 kHz. Inconsistency in $\Gamma_{\text{o}}$ in (b) is due to drifts in the mode matching between the optical pump and the transducer optical cavity over the longer averaging times required for downconversion spectrum measurements. Error bars indicate one standard deviation, and are dominated by statistical uncertainty in the fit parameters used to infer the measurement efficiency. Error bars in the $y$-direction are smaller than the data points. The black line is performance modeled using independently measured parameters. Transducers have nonzero quantum capacity for $N_{\text{add}}<1$ (light blue region). The quantum capacity of transducers in upconversion (c) and downconversion (d). Measurements presented in this work (maroon diamonds) are compared with other reported results (circles, labeled by first author)[25, 11, 10, 26, 27, 28, 13, 29, 20, 12] . The high throughput measured in downconversion in this work is enabled by our platform’s robustness to optical circulating power. Contours indicate $\mathcal{C}_{\text{ub}}$ as a function of a transducers throughput and $N_{\text{add}}$, assuming $\eta\ll 1$ and frequency-independent $N_{\text{add}}(\omega)$.
Turning to downconversion performance, the first three terms of Eq. 2 reveal that, in striking contrast with upconversion, $N_{\text{add,m}}$ can be made quite small simply by increasing $\Gamma_{\text{o}}$, because the excess motional noise can be routed out the input port by increasing the low-noise optical pump.
If motion of the membrane were the only source of noise, like it is in upconversion, our device would demonstrate quite asymmetrical performance, favoring operation in downconversion with large $\Gamma_{\text{o}}$. However, when our device operates in downconversion, the microwave field also contributes to $N_{\text{add}}$ through $N_{\text{add,e\kern-0.70004pt/\kern-1.04996pto}}$. Assuming $\Gamma_{\text{T}}=\Gamma_{\text{o}}+\Gamma_{\text{e}}$ to combine the last three terms of Eq. 2, we have
$N_{\text{add}}^{\text{down}}(\omega_{\text{m}})=\frac{n_{\text{th}}\gamma_{\text{m}}}{\Gamma_{\text{o}}}+n_{\text{om}}+\frac{\bar{n}_{\text{e}}\Gamma_{\text{o}}}{\Gamma_{\text{e}}}+\frac{n_{\text{min,e}}\Gamma_{\text{e}}}{\Gamma_{\text{o}}}.$ (4)
Again parameterizing $\bar{n}_{\text{e}}=a_{\text{e}}\Gamma_{\text{e}}+b_{\text{e}}$ as above, $N_{\text{add}}^{\text{down}}$ contains terms proportional to $\Gamma_{\text{o}}$ and $\Gamma_{\text{o}}/\Gamma_{\text{e}}$, so occupation of the microwave circuit prevents arbitrary improvement in downconversion performance with increasing $\Gamma_{\text{o}}$. Additionally, the last term in Eq. 4 indicates that finite sideband resolution contributing to nonzero $n_{\text{min,e}}$ limits the ratio $\Gamma_{\text{e}}/\Gamma_{\text{o}}$, causing increased noise at high microwave drive strength. Optimal downconversion performance is then achieved by tuning $\Gamma_{\text{o}}$ to balance $n_{\text{th}}\gamma_{\text{m}}$ against $a_{\text{e}}$, and $\Gamma_{\text{e}}/\Gamma_{\text{o}}$ to balance $n_{\text{min,e}}$ against $b_{\text{e}}$ (see Fig. 3(b) below).
.3 Noise and throughput in transducers
In general, there is a trade-off between the added noise of a transducer, and the rate of successful signal transmission as determined by the transduction efficiency and bandwidth. In platforms that pulse their operation to avoid additional unwanted noise, the reduced duty cycle can further limit this communication rate. Here, we identify a transducer’s throughput $\Theta$ [13, 12] quantitatively with the efficiency-bandwidth-duty-cycle product $\Theta=\eta BD$, as a way to characterize the total rate of a transducer. Roughly, it gives the expected rate that single photon signals are detected at the output of a transducer, when sent at a repetition rate of $B$.
To experimentally demonstrate the competition between throughput and noise in our device, in Figs. 3(a) and (b) we display results of varying $\Gamma_{\text{e}}$ while holding $\Gamma_{\text{o}}$ roughly fixed at $\Gamma_{\text{o}}=2\pi\cdot 11$ kHz, and quantify $N_{\text{add}}(\omega_{\text{m}})$ and $\Theta$. The data points corresponding to the measurements shown in Figs. 1 and 2 are marked with hearts. The black lines show the predicted performance of our transducer using independently calibrated measurements and modified versions of Eqs. 1 and 2 that account for lossy cavities and finite sideband resolution (Methods). These model curves illustrate the turnaround expected by increasing the throughput (by increasing $\Gamma_{\text{e}}$) from inspecting Eqs. 1 and 2. In upconversion, we see agreement between this model and our measurements, and in downconversion, we see qualitative agreement that reproduces the general trend of our measurement.
In Figs. 3(c) and (d), we put the respective performance measurements in upconversion and downconversion in the context of other results in the field, using a format similar to that first used in Ref. [12]. We see that the throughputs we achieve at the noise levels we measure are consistent with the best achieved in the field in upconversion, although unlike the device in Ref. [12], the turnaround due to thermomechanical motion at low throughputs prevents our current device from achieving quantum operation. Relative to our previous work [13], this turnaround occurs at higher throughput because of a $\gamma_{\text{m}}$ that is over an order-of-magnitude higher, as measured electromechanically by ringdown of the transduction mode. In downconversion, our platform achieves over three orders of magnitude higher throughput, at the cost of one order of magnitude increase in $N_{\text{add}}$, compared with the leading downconversion demonstration in the field [11].
.4 A rate for quantum transduction
For practical applications of a quantum transducer, improvements to both noise and throughput performance are needed in the field, and both metrics are important for the overall quantum information transfer rate. For example, a noisy quantum channel may necessitate the use of entanglement distillation protocols, consuming some quantity of low-fidelity entanglement in order to produce fewer entangled pairs of higher fidelity [30, 31]. A higher-throughput transducer would better enable distillation protocols by making a greater quantity of entangled pairs available within the lifetime of the quantum memories, while a lower-throughput and lower-noise transducer may be able to achieve a comparable total rate.
A channel’s quantum capacity gives the theoretical limit to the rate at which quantum information may be transmitted through it, and is therefore a possible metric to use in evaluating a transducer’s performance. In the context of general-capability quantum nodes linked by noisy and lossy channels, using local operations and classical communication (LOCC) to process transmitted quantum information (e.g. using a heralding protocol) is likely important, as they are necessary for recovering nonzero rates when losses exceed 50% and one-way deterministic state transfer becomes impossible [32]. When LOCC operations are allowed, the upper bound to the rate is referred to as the two-way classically assisted quantum capacity $\boldsymbol{c}$, which quantifies the ability of the channel to transmit one qubit of information, or distribute one ebit of entanglement, on average per attempted use of the channel [33, 34].
To calculate a rate of quantum information transfer possible for a realistic channel with added noise, limited efficiency, and finite bandwidth, we can integrate known upper bounds $\boldsymbol{c}_{\text{ub}}$ on the quantum capacity $\boldsymbol{c}$ for the thermal loss channel [34, 35], yielding an integrated capacity upper bound $\mathcal{C}_{\text{ub}}=\int\boldsymbol{c}_{\text{ub}}\text{d}{\omega}$. Assuming a frequency-independent $N_{\text{add}}$, which applies to our device’s operation in upconversion, we find that $\mathcal{C}_{\text{ub}}$ is approximately linear in the throughput $\Theta$ for small efficiencies, and is given by
$\mathcal{C}_{\text{ub}}(N_{\text{add}},\Theta)\approx\frac{\pi\;\Theta}{\ln(2)}\bigl{(}1-N_{\text{add}}+N_{\text{add}}\ln(N_{\text{add}})\bigr{)},$ (5)
for $N_{\text{add}}<1$ (see Methods for derivation). Though this approximation is pessimistic for channels with high efficiency, it remains accurate to within 17% for $\eta=0.5$, and to within 3% for $\eta=0.1$.
In the small-$\eta$ limit, that $\mathcal{C}_{\text{ub}}$ depends only on $\Theta$ and $N_{\text{add}}$ gives theoretical motivation to the choice of axes to summarize device performance in Ref. [12], and in Fig. 3. (For clarity, we here point out that, for example, the viability of deterministic state transfer protocols, which benefit from high efficiency, is not captured by $\Theta$ alone.) We can therefore overlay contours of $\mathcal{C}_{\text{ub}}$ onto Figs. 3(c) and (d), and gain an intuition for the relative importance of throughput and added noise in the quantum-enabled regime of $N_{\text{add}}<1$. The contours make it apparent that although reducing $N_{\text{add}}$ gives very sharp improvement while $N_{\text{add}}\approx 1$, further reductions in noise yield diminishing returns. Meanwhile, as $\mathcal{C}_{\text{ub}}$ is linear in $B$ and $D$, and at least linear in $\eta$, improving a channel’s throughput is always beneficial.
.5 Conclusions
Reasonable improvements to our device design would allow quantum-enabled noise performance. Refinement of the engineered mechanical isolation of our pinned membrane would suppress the stiff modes in the mechanical spectrum and address the elevated mechanical intrinsic loss rate measured in this device (Methods). Recovering the mechanical coherence we have measured in previous generations of this device, combined with parallel improvement to the electromechanical design to increase the electromechanical coupling rate [36], would yield high-throughput quantum-enabled performance in both upconversion and downconversion. With transduction demonstrations improving in both throughput and added noise, we are rapidly approaching transducer integrated quantum capacities at the kHz-scale, at which point transducers would support, on average, the entanglement of one pair of superconducting qubits within every 1 ms physical lifetime.